1. If average return in the stock market = 7.5% and average return on Treasury bills as a measure of risk free return = 1.5%, and beta of company Thor = 0.70.

And consider the following data on the company Thor.

SPS = Sales per share 5.70

EPS = Earnings per share 2.75

DPS = Dividends per share 1.25 (this is last year dividend or D0 )

BV = Book value per share 5.10

NPM = Net profit margin 7.2%

ROE = Return on equity 18.5%

MP = Market price per share 26.29

Normal growth in EPS of 3.5%.

Super growth rate of 8.5% for 12 years.

1.A. Calculate the Required Return on company Thor using the capital asset pricing model CAPM.

1.B.   Calculate the fair value based on the normal growth in earnings.

1.C. Compare your answer in part 1B and the data on the market price and state if you buy or sell.

Why?

1.D. Calculate the value of company Thor based on the super growth model.

1.E. Based on your answer in 1D, would you buy or sell?

Why?
Answer #1

1A. CAPM formula is as below

ERi​=Rf​+βi​(ERm​−Rf​); where:

ERi​=expected return of investment

Rf​=risk-free rate = 1.5%

βi​=beta of the investment = 0.70

(ERm​−Rf​)=market risk premium​ = (7.5% - 1.5%) = 6%

ERi = 1.5% + 0.70*6% = 1.5% + 4.2% = 5.7%

1B. P=d1 / (r-g), where P is the price, D1 is the dividend at year 1, r is req. rate of return, g is growth rate

d1 = DPS(1+g) = 1.25*(1+3.5%) = 1.29375

P=1.29375/(.057-.035)=58.8068 ; r came from Q 1A .

1C. yes, i would buy as the fair value is much higher than the market price and i am getting the stock cheaper.